OpenStudy (anonymous):

solve algebraically for x : 3(x+1)-5x=12-(6x-7)

OpenStudy (anx):

there are many different way to asnwer this, you can graph both and see where they intersect, but the answer is 4 and im fixing to show proof gimme a sec.

OpenStudy (anonymous):

I don't see where you would need to graph anything o.o First, you will need to distribute to get rid of those parentheses: 3x+3-5x=12-6x+7 Then you can add all factors that are alike: 3-2x=19-6x Now you want to get all of your x's on one side, I advise the left side. Add 6x to both sides: 3-2x+6x=19-6x+6x 3+4x=19 You still want the x to be by itself, so subtract 3 from both sides: -3+3+4x=19-3 4x=16 Now to get the 4 separated from the x you have to divide: 4x/4=16/4 x=4

OpenStudy (anonymous):

thankyou sharedjacket

OpenStudy (anonymous):

and anx!!